3ab^2+5a^2b-1 when a=2 and b=-2
Answer by venugopalramana(3286)

F=3ab^2+5a^2b-1 when a=2 and b=-2
JUST SUBSTITUTE THE GIVEN VALUES FOR A AND B AND COMPUTE THE ANSWER AS SHOWN BELOW
F=3*2*(-2)^2+5(2^2)*(-2)-1=6*4+5*4*(-2)-1=24-40-1=-17